OpenStudy (anonymous):

Let L1 be the line passing through the points Q1=(1, −1, −1) and Q2=(−3, 5, −5) and let L2 be the line passing through the point P1=(1, −5, 2) with direction vector →d=[−2, −1, 1]T. Determine whether L1 and L2 intersect. If so, find the point of intersection Q.

ganeshie8 (ganeshie8):

Express the lines in parametric form, set the components equal to each other and solve the parameter values

OpenStudy (amistre64):

you could also define the plane that would contain the lines, and see if the other lines anchor point is in the same plane

OpenStudy (anonymous):

how do i make them both in parametric form?

ganeshie8 (ganeshie8):

do you know how to write their equations in vector form ?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

can i use a matrix to solve for the parameters?

ganeshie8 (ganeshie8):

sure, you can

OpenStudy (anonymous):

is there an easier way?

OpenStudy (anonymous):

will x,y,z all be equal on both lines?

OpenStudy (amistre64):

project them onto the xy plane and see where they cross ... then determine if they have the same z value for that points not sure if it 'easier'

ganeshie8 (ganeshie8):

solving system of 3 equations shouldn't be hard i guess ?

ganeshie8 (ganeshie8):

it is painful, but since you're doing linear algebra you should be good at solving system of linear equations :)

OpenStudy (amistre64):

q1=(1, −1) q2=(−3, 5) ---------- 4,-6, slope of -3/2 p1=(1, −5) with direction vector [−2, −1], slope of 1/2

OpenStudy (anonymous):

fml thanks you are really helpful

OpenStudy (anonymous):

amsitre64 i'm not really understanding the "planes" way of solving it

ganeshie8 (ganeshie8):

so are you able to figure out how to express the given line in vector form yet ?

ganeshie8 (ganeshie8):

*lines

OpenStudy (amistre64):

cross the direction vectors, anchor it to one of the points, and see if the other points fit in the plane .... lines only cross if they can fit in the same plane

OpenStudy (anonymous):

i think i can yeah

OpenStudy (amistre64):

 if all the points given are in the same plane ... then the lines cross but its not really going to tell us where they cross at